60kg had collision with40kgman with the force of60NWHAT WILLBE THE FORCE EXPERIENCED BY THE 40KG MAN

Grade:12th pass

1 Answers

John Sharon Sandeep
6 years ago
Now, when the 60kg object contacts the 40kg man, the force remains 60N, but now 60N is not only applied on the object but also the man.....So total mass on which force is acting=60kg+40kg=100kg. Therefore now we need to find the new rate at which the man(and object) is accelerating before finding the force acting on him....So a=F/m=60N/100kg=0.6ms-1.☺ therefore force acting on the man seperately=F=ma=40kg×0.6ms-1= 24N😊😊